You found the following stock quote for Gigantus Corporation in today’s newspaper. What was the stock...

  1. You found the following stock quote for Gigantus Corporation in today’s newspaper. What was the stock selling for yesterday? What was the stock selling for on January 1?

Company: Symbol: Vol: Close: Daily Chg: %chg: YTD %chg: High: 52 week low: %chg:

Gigantus GIG 12,805,325 49.51 0.78    1.60% -3.20% 62.81 45.93   6.50%
Answer #1

The following information is given

Company Symbol Vol Close Daily Chg % Chg YTD % Chg High 52 week Low
Gigantus GIG 12805325 49.51 0.78 1.60% -3.20% 62.81 45.93

One can observe that today's closing price was 49.51 and daily change was 0.78 . it means that stock has risen by 0.78 from it's previous close

Yesterday's selling price = 49.51-0.78 = 48.73

Now YTD( year to date change is given) i.e -3.2%. It means that stock has fallen by 3.2% from the price it was traded one year ago ( Date is not mentioned in question hence i am assuming that 1st January refers to one year before date)

Assuming price 1 year before be X

X - 0.032(X) = 49.51

X-0.032X = 49.51

0.968X = 49.51

X =51.1467 i.e 51.15

Hence price on January 1 was 51.15